# Neural Differential Algebraic Equations

### v1

@@ 6       $x^{*}\longleftarrow\textnormal{{ODESolve}}\left(f,\{x,y^{*},u\};\theta_{f}\right)$
8 while not converged do
@@ 12             DAESolve($f,h,\{x^{(k-1)},y^{(k-1)},u^{(k-1)}\}$) )
13
      /* Evaluate loss */
14       $\mathcal{L}\longleftarrow$ Criterion ( $X,\hat{X}$ )
15       if converged then
@@ 14       $\mathcal{L}\longleftarrow$ Criterion ( $X,\hat{X}$ )
19             obtain gradients;
20             update model weights $\{\theta_{f},\theta_{h}\}$;
21
@@
Algorithm 1 Neural DAE training
IV NUMERICAL CASE STUDIES
All experiments were performed using the NeuroMANCER Scientific Machine Learning (SciML) package [31] built on top of PyTorch [32] [^1] Repository containing all training and visualization scripts will be open-sourced and publicly available by the time of publication.. Each presented model was trained using an Intel I9-9980 CPU laptop with a wall-clock training time on the order of 5 minutes.
IV-A Tank-Manifold Property Inference
IV-A1 Problem Description
A manifold is a physical realization of a conservation relationship that manifests as an algebraic constraint in a DAE. Figure  2 depicts a pair of tanks fed by a manifold. A pseudo-infinite reservoir supplies the pump. At the manifold, the inflow is exactly equal to the sum of the outflows; that is, $y_{\text{in}}=y_{1}+y_{2}$, where $y$ represents volumetric flow rates and the subscripts $1$ and $2$ denote the two exit streams of the manifold. In this example, the two tanks have different area-height profiles but have a common datum such that the column heights are identical. As a fluid is pumped into the two tanks, the flow rates in the two streams varies to ensure (i) the conservation relationship of the manifold and (ii) that each of the tanks have identical pressure head, i.e $x_{1}=x_{2}$, where $x$ is the height of the fluid in the respective tank. The full DAE that describes the evolution of this system is given as:
@@ IV-A1 Problem Description
where $\phi_{1}$ and $\phi_{2}$ denote the area-height profiles for the two tanks. The solution of Eq.  11 can be obtained with many off-the-shelf standard scientific computing packages. For this study, the area-height profiles are set to $\phi_{1}(x)=3.0$ and $\phi_{2}(x)=\sqrt{x}+0.1$ respectively. The goal of the learning task is to recover $\phi_{2}(x)$ given observations of tank heights $x$ and pump volumetric flow rates $y$. Note that if the task were to recover both $\phi_{1}(x)$ and $\phi_{2}(x)$ simultaneously, the problem would be ill-posed in the sense that there does not exist a unique set of tank profiles that recovers the behavior seen in observations.
@@
IV-A2 Model Construction
In the system described by Eq.  11, the differential states are column heights $x_{1}$ and $x_{2}$ and the algebraic states are the volumetric flows $y_{1}$ and $y_{2}$. The input flow is treated as an exogenous control input, i.e., $u=y_{\text{in}}$.
We aim to approximate the area-height profile of Tank #1 as well as construct a surrogate for the behavior of the manifold such that (i) fluid is conserved and (ii) the fluid column height constraint is satisfied. There are two neural networks that we seek to train: $NN_{1}:\mathbb{R}^{1}\rightarrow\mathbb{R}^{1}$, an approximation of the sought area-height profile $\phi_{1}$, and $NN_{2}:\mathbb{R}^{4}\rightarrow(0,1)$, the surrogate of the behavior of the manifold. Each are implemented as a feed-forward neural network with one hidden layer of 5 nodes, each with sigmoid activation functions. The output layer of $NN_{1}$ is linear while the output layer of $NN_{2}$ is sigmoidal, constraining the output to $(0,1)$.
IV-A3 Optimization Problem
A time series dataset is constructed by simulation of the system given in Eq.  11 regularly sampled on the interval $t\in[0,500]$ with $\Delta t=1.0$ and with a constant inflow of $y_{\text{in}}=u=0.5$:
$\hat{X}=\{(\hat{x}^{(0)},\hat{y}^{(0)}),...,(\hat{x}^{(N)},\hat{y}^{(N)})\},$ (12)
where $N$ is the time series prediction horizon and the $\widehat{\cdot}$ denotes membership to the dataset. The optimization problem seeks the minimum of the differential state trajectory reconstruction loss and algebraic state reconstruction loss when predicted by our $\operatorname{DAESolve}$:
@@ IV-A3 Optimization Problem
Note that the mapping $h$ defines a convex combination at the manifold: the outflows $y_{1}$ and $y_{2}$ will sum to $u$ by construction.
For all examples contained in this work, the optimizer is Adam [30] with fixed learning rate of 0.001. Training is deemed ‘converged’ after 20,000 epochs or until an early stop is triggered after 20 epochs of increasing loss.
IV-A4 Results
@@
TABLE I: Mean Square error for the differential state and algebraic states for tanks-manifold-pump experiment.
@@
IV-B Tank Network Modeling
IV-B1 Problem Description
Figure 5 expands upon the tank-manifold problem of Fig. 2 by adding additional tanks, feedback mechanisms, and ‘closing the loop’ in the network flow. In this example, the pump flow rate depends on the fluid heights of Tank #1 and the reservoir. Tanks #1 and #2 have a height constraint (equivalent pressure head) and are fed by a common manifold. The outlets of tanks #1 and #3 depend on the square root of the fluid column height in the respective tanks. The full system can be described by the DAE:
@@ IV-B1 Problem Description
where $x$ denotes the fluid column height for tanks #1 through #4 (the reservoir), $p$ is a pump response function, $y$ represents volumetric flows, and $\alpha$ describe the discharge coefficients of the tank outlets. The differential states are $x=\{x_{1},x_{2},x_{3},x_{4}\}\in\mathbb{R}^{4}$ and the algebraic states are $y=\{y_{0},y_{1},y_{2},y_{3},y_{4}\}\in\mathbb{R}^{5}$.
In this example, several characteristics make the dynamics more complex. First, the system is closed, as there is no material transfer into or out of the system. Second, there is the potential for backflow through the manifold from Tank #2. Lastly, the feedback mechanism from the tank level transducers to the pump controller is such that oscillations can occur.
@@ IV-B2 Optimization Problem
A single-trajectory dataset $\hat{X}$ was constructed through sampling this simulation at $\Delta t=0.1$ for $t\in[0,20]$. The optimization problem definition is identical to that of Eq. 13. For this problem, it is reasonable to assume that the basic physics for a gravity-fed pipe network are known, which provides a starting point for constructing a model. However, discharge coefficients and quantifying feedback mechanisms are not as straightforward. Here, we attempt to approximate (i) the Level Controller (LC) dynamics for pump control with a neural network $\operatorname{NN}_{1}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{1}$, (ii) manifold dynamics with with a second neural network $\operatorname{NN}_{2}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{1}$, and (iii) the discharge coefficients $\alpha_{1}$ and $\alpha_{2}$. The mappings $f$ and $h$ are defined as:
@@
For this example, both $\operatorname{NN}_{1}$ and $\operatorname{NN}_{2}$ are simple multi-layer perceptrons each with a two hidden layers of size 30 and ReLU activations.
IV-B3 Results
Figure 6(a) details the time history of the four tank heights for both the ground truth data as well as the open-loop model rollout. At these process conditions, the pump delivers an oscillatory flow to the manifold, which our model is able to capture. Additionally, the tuned model is able to infer the values for the discharge coefficients: for the case of zero noise the recovered values are $\alpha_{1}=0.1027$ and $\alpha_{2}=0.1024$.
To demonstrate generalizability, the trained model is deployed on an unseen area-height tank profile; for this, we redefine $\phi_{1}(x_{1})=1.0$. In doing so, the system’s limit cycle is destroyed, and the true dynamics approach a steady state. Figure  6(b) demonstrates this extrapolation for the same time horizon. The reconstructed time series data matches the behavior of the unseen data: there exist damped oscillations that decay to a steady state value. Table-II shows the MSE for the neural DAE models trained for the tank network problem.
Noise tolerance is an important consideration in the evaluation of system identification tasks, especially in settings with nonlinearities and feedback mechanisms. For this example, we investigate noise tolerance by introducing Gaussian white noise. Figures  7(a) and  7(b) show similar training and extrapolation tasks for the Signal-to-Noise Ratio (SNR) of $20$ dB. Despite the significant noise, the learned dynamics are faithful to the underlying physics, as demonstrated in the extrapolation task, and result in an MSE of $027e-$ 02.
Figure 6: Tank Network problem : (a) Ground truth data vs Neural-DAE model prediction. (b) Extrapolation case: inferring behavior for unseen component parameters using Neural-DAE.
@@ IV-B3 Results
TABLE II: Mean Square error for the differential state and algebraic states for tanks network experiment.
@@
V CONCLUSIONS AND FUTURE WORKS
This work presents a framework for the data-driven modeling of Differential Algebraic Equations (DAEs). Specifically, we introduce a novel method called Neural Differential Algebraic Equations (NDAEs) that extends the notion of Universal Differential Equations to endogenize the evolution of algebraic relationships. Inspired by Lie-Trotter operator splitting, we achieve this by splitting the integration of system states into two sequential sub-tasks: first, updating algebraic states via a differentiable surrogate of an algebra solver, followed by a conventional temporal integration step for a system of Ordinary Differential Equations.
The effectiveness of these methods was demonstrated by (i) modeling networked dynamical systems with underlying constraints and conservation relationships and (ii) showing generalization to unseen process conditions or component variation. We demonstrate that the proposed NDAE method is robust to noise and possesses the extrapolation ability to (i) learn the behaviors of the DAE system components and their interaction physics and (ii) disambiguate between data trends and mechanistic relationships contained in the DAE system.
While the methods and example application problems were successful in this work, future research and development are needed in several areas. First, a central assumption of our method is that the underlying dynamics of the systems of interest are sufficiently smooth, i.e., not containing jump discontinuities or shocks, such that the underlying dynamics are non-stiff. As presented, our method leverages an explicit fractional-step approach to decoupling the evolution of the differential and algebraic states. To increase the ability of the method to handle stiff dynamics or discontinuities, moving to an implicit integration scheme may be warranted. Second, further analysis is needed to investigate the interplay between the modeling task and the index of the underlying DAE (for example, higher-order DAEs that are characteristic of design or parameter optimization problems). Lastly, we have made no claims about computational scalability nor partial observability, both of which are critical for modeling real-world complex systems with hundreds of degrees of freedom. We reserve these topics for future work.

### v2

@@ 6       $x^{*}\longleftarrow\textnormal{{ODESolve}}\left(f,\{x,y^{*},u\};\theta_{f}\right)$
8 while $\mathcal{L}(\Theta)>\epsilon$ do
9
@@ 13             DAESolve($f,h,\{x^{(k-1)},y^{(k-1)},u^{(k-1)}\}$) )
14
15
      /* Evaluate loss via Eq. (11) */
16       $\mathcal{L}(\Theta)\longleftarrow$ Criterion ( $X,\hat{X}$ )
17
18      if $\mathcal{L}(\Theta)\leq\epsilon$ then
@@ 16       $\mathcal{L}(\Theta)\longleftarrow$ Criterion ( $X,\hat{X}$ )
            /* Backward pass */
23             $\nabla\mathcal{L}(\Theta)\longleftarrow\frac{\partial\mathcal{L}(\Theta)}{\partial\Theta}$
            /* Update model parameters */
24             $\Theta\longleftarrow\Theta-\eta\nabla\mathcal{L}(\Theta)$
25
Algorithm 1 Operator splitting-based DAE training
V NUMERICAL CASE STUDIES
All experiments were performed using the NeuroMANCER Scientific Machine Learning (SciML) package [36] built on top of PyTorch [37]. The open-source implementation of the presented method and associated numerical examples are available at GitHub [^2] https://github.com/pnnl/NeuralDAEs. Each presented model was trained using an Intel I9-9980 CPU laptop with a wall-clock training time on the order of 5 minutes.
V-A Tank-Manifold Property Inference
V-A1 Problem Description
A manifold is a physical realization of a conservation relationship that manifests as an algebraic constraint in a DAE. Figure 2 depicts a pair of tanks fed by a manifold. A pseudo-infinite reservoir supplies the pump. At the manifold, the inflow is exactly equal to the sum of the outflows; that is, $y_{\text{in}}=y_{1}+y_{2}$, where $y$ represents volumetric flow rates and the subscripts $1$ and $2$ denote the two exit streams of the manifold. In this example, the two tanks have different area-height profiles but have a common datum such that the column heights are identical. As a fluid is pumped into the two tanks, the flow rates in the two streams varies to ensure (i) the conservation relationship of the manifold and (ii) that each of the tanks have identical pressure head, i.e $x_{1}=x_{2}$, where $x$ is the height of the fluid in the respective tank. The full DAE that describes the evolution of this system is given as:
@@ V-A1 Problem Description
where $\phi_{1}$ and $\phi_{2}$ denote the area-height profiles for the two tanks. The solution of Eq. (12) can be obtained with many off-the-shelf standard scientific computing packages. For this study, the area-height profiles are set to $\phi_{1}(x)=3.0$ and $\phi_{2}(x)=\sqrt{x}+0.1$ respectively. The goal of the learning task is to recover $\phi_{2}(x)$ given observations of tank heights $x$ and pump volumetric flow rates $y$. Note that if the task were to recover both $\phi_{1}(x)$ and $\phi_{2}(x)$ simultaneously, the problem would be ill-posed in the sense that there does not exist a unique set of tank profiles that recovers the behavior seen in observations.
@@
V-A2 Model Construction
In the system described by Eq. (12), the differential states are column heights $x_{1}$ and $x_{2}$ and the algebraic states are the volumetric flows $y_{1}$ and $y_{2}$. The input flow is treated as an exogenous control input, i.e., $u=y_{\text{in}}$. We aim to approximate the area-height profile of Tank #1 as well as construct a surrogate for the behavior of the manifold such that (i) fluid is conserved and (ii) the fluid column height constraint is satisfied. There are two neural networks that we seek to train: $NN_{1}:\mathbb{R}^{1}\rightarrow\mathbb{R}^{1}$, an approximation of the sought area-height profile $\phi_{1}$, and $NN_{2}:\mathbb{R}^{4}\rightarrow(0,1)$, the surrogate of the behavior of the manifold. Each is implemented as a feed-forward neural network with one hidden layer of 5 nodes with sigmoid activation functions. The output layer of $NN_{1}$ is linear while the output layer of $NN_{2}$ is sigmoidal, constraining the output to $(0,1)$.
V-A3 Optimization Problem
A time series dataset $\hat{X}$ is constructed by simulation of the system given in Eq. (12) regularly sampled on the interval $t\in[0,500]$ with $\Delta t=1.0$ and with a constant inflow of $y_{\text{in}}=u=0.5$. The optimization problem seeks the minimum of the differential state trajectory reconstruction loss and algebraic state reconstruction loss when predicted by our $\operatorname{DAESolve}$:
@@ V-A3 Optimization Problem
Note that the mapping $h$ defines a convex combination at the manifold: the outflows $y_{1}$ and $y_{2}$ will sum to $u$ by construction. For all examples contained in this work, the optimizer is Adam [35] with a fixed learning rate of 0.001. Training is deemed ‘converged’ after 20,000 epochs or until an early stop is triggered after 20 epochs of increasing loss.
V-A4 Results
@@
TABLE I: Mean Square error for the differential state and algebraic states for the tanks-manifold-pump experiment.
@@
V-B Tank Network Modeling
V-B1 Problem Description
Figure 5 expands upon the tank-manifold problem of Fig. 2 by adding additional tanks, feedback mechanisms, and ‘closing the loop’ in the network flow. In this example, the pump flow rate depends on the fluid heights of Tank #1 and the reservoir. Tanks #1 and #2 have a height constraint (equivalent pressure head) and are fed by a common manifold. The outlets of tanks #1 and #3 depend on the square root of the fluid column height in the respective tanks. The full system is described by DAEs:
@@ V-B1 Problem Description
where $x$ denotes the fluid column height for tanks #1 through #4 (the reservoir), $p$ is a pump response function, $y$ represents volumetric flows, and $\alpha$ describe the discharge coefficients of the tank outlets. The differential states are $x=\{x_{1},x_{2},x_{3},x_{4}\}\in\mathbb{R}^{4}$ and the algebraic states are $y=\{y_{0},y_{1},y_{2},y_{3},y_{4}\}\in\mathbb{R}^{5}$. In this example, several characteristics make the dynamics more complex. First, the system is closed, as there is no material transfer into or out of the system. Second, there is the potential for backflow through the manifold from Tank #2. Lastly, the feedback mechanism from the tank level transducers to the pump controller is such that oscillations can occur.
@@ V-B2 Optimization Problem
A single-trajectory dataset $\hat{X}$ was constructed through sampling this simulation at $\Delta t=0.1$ for $t\in[0,20]$. The optimization problem definition is identical to that of Eq. (13). For this problem, it is reasonable to assume that the basic physics for a gravity-fed pipe network is known, which provides a starting point for constructing a model. However, discharge coefficients and quantifying feedback mechanisms are not as straightforward. Here, we attempt to approximate (i) the Level Controller (LC) dynamics for pump control with a neural network $\operatorname{NN}_{1}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{1}$, (ii) manifold dynamics with with a second neural network $\operatorname{NN}_{2}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{1}$, and (iii) the discharge coefficients $\alpha_{1}$ and $\alpha_{2}$. The mappings $f$ and $h$ are defined as:
@@
For this example, both $\operatorname{NN}_{1}$ and $\operatorname{NN}_{2}$ are simple multi-layer perceptrons, each with two hidden layers of size 30 and ReLU activations.
V-B3 Results
Figure 6(a) details the time history of the four tank heights for both the ground truth data and the open-loop model rollout. At these process conditions, the pump delivers an oscillatory flow to the manifold, which our model is able to capture. Additionally, the tuned model can infer the values for the discharge coefficients: for the case of zero noise, the recovered values are $\alpha_{1}=0.1027$ and $\alpha_{2}=0.1024$. To demonstrate generalizability, the trained model is deployed on an unseen area-height tank profile; for this, we redefine $\phi_{1}(x_{1})=1.0$. In doing so, the system’s limit cycle is destroyed, and the true dynamics approach a steady state. Figure 6(b) demonstrates this extrapolation for the same time horizon. The reconstructed time series data matches the behavior of the unseen data: there exist damped oscillations that decay to a steady state value. Table-II shows the MSE for the neural DAE models trained for the tank network problem. Noise tolerance is an important consideration in the evaluation of system identification tasks, especially in settings with nonlinearities and feedback mechanisms. For this example, we investigate noise tolerance by introducing Gaussian white noise. Figures 7(a) and 7(b) show similar training and extrapolation tasks for the Signal-to-Noise Ratio (SNR) of $20$ dB. Despite the significant noise, the learned dynamics are faithful to the underlying physics, as demonstrated in the extrapolation task, and result in an MSE of $027e-$ 02.
Figure 6: Tank Network problem: (a) Ground truth data vs NDAE model prediction. (b) Extrapolation case: inferring behavior for unseen component parameters using NDAE.
@@ V-B3 Results
TABLE II: Mean Square error for the differential state and algebraic states for the tanks network experiment.
@@
VI CONCLUSIONS AND FUTURE WORKS
This work presents a framework for the data-driven modeling of Differential Algebraic Equations (DAE). Specifically, we introduce a novel Operator Splitting (OS) integration method for Neural Differential Algebraic Equations (NDAE). Inspired by Lie-Trotter operator splitting, we achieve this by splitting the integration of system states into two sequential sub-tasks: first, updating algebraic states via a differentiable surrogate of an algebra solver, followed by a conventional temporal integration step for a system of Ordinary Differential Equations.
The effectiveness of the proposed method is demonstrated by (i) modeling networked dynamical systems with underlying constraints and conservation relationships and (ii) showing generalization to unseen process conditions or component variation. We demonstrate that the proposed OS-based NDAE method is robust to noise and possesses the extrapolation ability to (i) learn the behaviors of the DAE system components and their interaction physics and (ii) disambiguate between data trends and mechanistic relationships contained in the DAE system.
While the methods and example application problems were successful in this work, future research and development are needed in several areas. First, a central assumption of our method is that the underlying dynamics of the systems of interest are sufficiently smooth, i.e., not containing jump discontinuities or shocks, such that the underlying dynamics are non-stiff. As presented, our method leverages an explicit operator splitting method to decouple the evolution of the differential and algebraic states. To increase the ability of the method to handle stiff dynamics or discontinuities, moving to an implicit integration scheme may be warranted. Second, further analysis is needed to investigate the interplay between the modeling task and the index of the underlying DAE (for example, higher-order DAEs that are characteristic of design or parameter optimization problems). Lastly, we have made no claims about computational scalability or partial observability, both of which are critical for modeling real-world complex systems with hundreds of degrees of freedom. We reserve these topics for future work.
@@ ACKNOWLEDGMENTS
This research was also supported by the Ralph O’Connor Sustainable Energy Institute at Johns Hopkins University.